OpenStudy (anonymous):

$5,000 is invested in two different accounts yielding 3% and 3.5% interest. The interest earned on the two accounts is $155. Let x and y represent the two amounts invested. Complete the equation. = 155 A.3x+3.5y B.0.3x+0.35y C.0.03x+0.035y i think its b

OpenStudy (anonymous):

x+y=5000 => y=5000-x 0.03x+0.035(5000-x)=155 0.03x+175-0.035x=155 0.005x=20 x=4000 So, C

OpenStudy (anonymous):

do you follow me?

OpenStudy (anonymous):

if you explain it then ill get it XP so thanks

OpenStudy (e.mccormick):

When you convert from % to decimal, you go over 2. 100% = 1, 50% = .5, 1% = .01, etc. 

OpenStudy (e.mccormick):

So when you convert the % to decimals, it must go over 2 spots, making any answer that does not do that an impossible choice.

OpenStudy (e.mccormick):

What ryanvarghese12790 did was see what would give a valid answer by testing it mathematically. What I did was look at the rules of math and eliminate impossible choices logically. Both are valid ways to get the right answer.